### v2

# Analog Computing for Signal Processing and Communications – Part II: Toward Gigantic MIMO Beamforming

Abstract
Analog-domain operations offer a promising solution to accelerating signal processing and enabling future multiple-input multiple-output (MIMO) communications with thousands of antennas. In Part I of this paper, we have introduced a microwave linear analog computer (MiLAC) as an analog computer that processes microwave signals linearly, demonstrating its potential to reduce the computational complexity of specific signal processing tasks. In Part II of this paper, we extend these benefits to wireless communications, showcasing how MiLAC enables gigantic MIMO beamforming entirely in the analog domain. MiLAC-aided beamforming enables the maximum flexibility and performance of digital beamforming, while significantly reducing hardware costs by minimizing the number of radio-frequency (RF) chains and only relying on low-resolution analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). In addition, it eliminates per-symbol operations by completely avoiding digital-domain processing and remarkably reduces the computational complexity of zero-forcing (ZF), which scales quadratically with the number of antennas instead of cubically. It also processes signals with fixed matrices, e.g., the discrete Fourier transform (DFT), directly in the analog domain. Numerical results show that it can perform ZF and DFT with a computational complexity reduction of up to $1.5\times 10^{4}$ and $4.0\times 10^{7}$ times, respectively, compared to digital beamforming.
Index Terms:
Analog beamforming, analog computing, massive multiple-input multiple-output (MIMO), minimum mean square error (MMSE), zero-forcing (ZF).
I Introduction
In Part I of this paper, we have reviewed the growing interest in analog computing, driven by advancements in several technologies such as optical systems, metamaterials, and resistive memory arrays. Digital computing, despite its dominance, faces fundamental limitations in speed and power consumption, mainly due to the reliance on analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). Analog computing addresses these challenges by enabling energy-efficient and highly parallelized computations at the speed of light. While Part I has focused on how analog computing can be leveraged for signal processing operations, Part II explores how these principles can be applied to achieve efficient and flexible beamforming in the analog domain, paving the way for future wireless communications.
As we are witnessing a revival of interest in analog operations within the field of signal processing, a similar shift from digital to analog has emerged in wireless communications. This shift is driven by the growing number of antennas needed to support higher data rates and simultaneously serve a larger number of users. While massive multiple-input multiple-output (MIMO) systems typically deploy 64 or more antennas [2, 3], this number could reach several thousands in ultra-massive or gigantic MIMO systems, which have been recently proposed to meet the stringent requirements of 6G networks [4]. With such a large number of antennas, digital beamforming faces two critical challenges in terms of hardware and computational complexity. First, digital beamforming typically requires a dedicated radio-frequency (RF) chain per antenna element, each including ADCs/DACs and mixers, which are costly and power-hungry components. Second, as the number of antennas increases, the data volumes to be processed in real-time grow accordingly. For example, to precode the transmitted symbol vector at the transmitter (or combine the received symbol vector at the receiver) a matrix-vector product is required on a per-symbol basis. These two challenges result in high hardware costs and computational complexity when digital beamforming is applied in massive or gigantic MIMO systems. For these reasons, several alternatives to digital beamforming have been explored to enable beamforming operations fully or partially in the analog domain.
To address the high hardware cost associated with digital beamforming, alternative strategies such as analog beamforming and hybrid analog-digital beamforming have been proposed. Analog beamforming steers the beam pattern by solely requiring a single RF chain [5]. This is achieved by connecting the RF chain to the antennas via phase shifters or time delay elements, which impose a constant modulus constraint on the beamforming weights. Thus, while analog beamforming offers low hardware complexity, it is characterized by limited flexibility. To balance hardware cost and flexibility by bridging analog and digital beamforming, hybrid analog-digital beamforming techniques have been widely investigated [5, 6, 7, 8, 9]. In hybrid beamforming, the beamforming process is divided into two stages. First, the transmitted symbols are digitally processed in the baseband domain with a low-dimensional digital beamforming matrix, and fed to a reduced number of RF chains. Second, the RF chains are connected to the transmitting antennas through a network of phase shifters, operating analog beamforming in the RF domain. The reversed process is applied at the receiver.
Another method to efficiently perform beamforming in the analog domain is given by reconfigurable intelligent surface (RIS) [10, 11]. This technology has recently emerged enabling the control of the wireless channel through surfaces made of multiple passive elements with reconfigurable scattering properties. Beamforming in the analog domain can be achieved through RIS by deploying a RIS in close proximity to an active transceiving device equipped with a reduced number of antennas. In this way, the effective radiation pattern of the active device can be optimized by reconfiguring the RIS elements [12, 13, 14]. Two extensions of this technology have been proposed to achieve additional flexibility over conventional RIS. First, beyond diagonal RIS (BD-RIS) include more general RIS architectures, characterized by a scattering matrix not restricted to be diagonal [15, 16, 17]. By deploying a BD-RIS close to an active device, its additional flexibility over conventional RIS in manipulating the radiation pattern can enable even higher performance [18]. Second, stacked intelligent metasurface (SIM) technology considers multiple stacked transmissive RISs deployed close to the active device [19, 20, 21]. The additional flexibility provided by the multiple layers has shown increased performance over just deploying a single conventional RIS.
In the reviewed beamforming strategies operating in the analog domain, the number of RF chains is commonly lower bounded by the number of symbols, or streams, transmitted in parallel. Nevertheless, other beamforming strategies have been developed to also perform symbol modulation in the analog domain, allowing the transmission of multiple symbols with a single RF chain. First, in electronically steerable parasitic array radiators (ESPARs), the symbols are modulated and the radiation pattern reconfigured by adjusting the mutual coupling between one active antenna element and multiple “parasitic” elements having tunable loads [22, 23, 24]. Second, in load modulated arrays (LMAs), a single RF source is connected to multiple antennas through tunable impedance components, which are reconfigured to achieve symbol modulation [25, 26]. Third, symbol modulation can be performed in the analog domain by employing a RIS deployed close to an active antenna transmitting a carrier signal [27, 28]. These three strategies further highlight the potential of leveraging analog-domain beamforming to reduce the hardware cost of transceiver devices.
In addition to alleviating the hardware cost, analog-domain strategies have been also explored to reduce the computational complexity of future wireless networks, by performing specific mathematical operations in the analog domain. To this end, in-memory computing has been proposed to accelerate ridge regression with applications to massive MIMO [29]. SIM has been used to compute the discrete Fourier transform (DFT) of the incident signal in the analog domain, which is particularly beneficial for direction-of-arrival (DoA) estimation [30]. Moreover, over-the-air computing emerged as a form of analog computing that leverages the superposition principle characterizing the electromagnetic (EM) propagation over the wireless channels. Over-the-air computing allows efficient computation of the so-called nomographic functions [31, 32, 33], with impactful applications on federated learning [34].
Numerous technologies have been proposed to reduce the hardware cost of future wireless transceivers through analog beamforming [5]-[28]. However, operating in the analog domain typically decreases the flexibility compared to digital beamforming, with a consequent reduction in performance. Thus, whether it is possible to maintain maximum performance while simultaneously minimizing hardware cost and computational complexity through analog computing [29]-[34] remains an open question. In Part II of this paper, we address this gap by introducing a novel beamforming strategy enabled by microwave linear analog computers (MiLACs), denoted as MiLAC-aided beamforming. This beamforming strategy maintains the same flexibility as digital beamforming while operating entirely in the analog domain. Because of its fully analog nature, it minimizes the number of RF chains and relies only on low-resolution ADCs/DACs, reducing the hardware cost. By leveraging the computational capabilities of MiLACs, it also minimizes the computational complexity required for beamforming. More in detail, the contributions of Part II of this paper are as follows.
First, we demonstrate that a MiLAC can efficiently compute five mathematical operations of practical interest in MIMO communications, such as least squares (LS) and matched filtering (MF). To this end, we first show how these five operations can be seen as special cases of the linear minimum mean square error (LMMSE) estimator for linear observation processes, and then discuss how the microwave network of the MiLAC needs to be reconfigured to compute them. These operations can be computed with a MiLAC with significantly reduced computational complexity compared to conventional digital computing.
Second, we propose MiLAC-aided beamforming as a new beamforming strategy enabled by MiLAC scalable to gigantic MIMO systems. MiLAC-aided beamforming can perform precoding at the transmitter and combining at the receiver with arbitrary matrices, i.e., with the same flexibility as digital beamforming. While having the maximum flexibility, MiLAC-aided beamforming allows us to perform specific operations in the analog domain, such as zero-forcing (ZF) precoding at the transmitter and ZF combining at the receiver, with important gains in terms of computational complexity. In addition, it allows to precode or combine a signal with a fixed matrix, such as the DFT matrix, entirely in the analog domain, requiring no digital operations.
Third, we compare MiLAC-aided beamforming with existing beamforming strategies, namely digital, analog, hybrid, RIS-aided, and SIM-aided beamforming, in terms of flexibility of the beamforming matrix, hardware complexity, and computational complexity. As a result of this comparison, we discuss five benefits of MiLAC-aided beamforming: 1) it enables the maximum flexibility, 2) it requires the minimum number of RF chains, 3) it only needs low-resolution ADCs/DACs, 4) it does not require any computation on a per-symbol basis, and 5) it requires the minimum number of computations per coherence block for ZF.
Fourth, we provide numerical results to assess the performance of MiLAC-aided beamforming in terms of sum rate and bit error rate (BER), confirming that it can achieve the same performance as digital beamforming. We also assess the benefits of MiLAC-aided beamforming over digital beamforming in terms of computational complexity, showing that it can perform ZF, MF, and the DFT with a gain in computational complexity of $1.5\times 10^{4}$, $2.0\times 10^{2}$, and $4.0\times 10^{7}$ times, respectively. Thus, in a gigantic MIMO system with thousands of antennas, MiLAC-aided beamforming achieves the same performance as digital ZF beamforming at just $1/15000$th of the computational cost. From a different perspective, a 4096-antenna MiLAC serving 4096 users via ZF beamforming requires the same computational complexity as a 256-antenna digital transmitter serving 256 users. This makes MiLAC-aided beamforming a game changer for gigantic MIMO, opening the door to radically new transceiver architectures that can massively scale the number of antennas and users.
Organization: In Section II, we briefly review the concept of MiLAC, as introduced in Part I of this paper. In Section III, we derive five special cases of the LMMSE estimator and show how they can be efficiently computed by a MiLAC. In Section IV, we propose MiLAC-aided beamforming as a novel beamforming strategy enabled by a MiLAC. In Section V, we compare MiLAC-aided beamforming with existing beamforming strategies. In Section VI, we provide numerical results to evaluate the performance and computational complexity of MiLAC-aided beamforming. Finally, Section VII concludes this paper.
Notation: Please, refer to the notation introduced in Section I of Part I of this paper.
Figure 1: Representation of a $P$-port MiLAC with input on $N$ ports, where $P=N+M$.
II Microwave Linear Analog Computer
In this section, we briefly review the concept of MiLAC, defined as an analog computer that linearly processes microwave signals, as introduced in Part I of this paper. In general, a MiLAC can be seen as a $P$-port reconfigurable microwave network, which can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$. We assume this microwave network to be made of $P^{2}$ tunable admittance components interconnecting each port to ground and to all the other ports, such that the microwave network can be reconfigured to assume any arbitrary admittance matrix $\mathbf{Y}$. The admittance component interconnecting port $k$ to ground is denoted as $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, while the admittance component interconnecting port $k$ to port $i$ is denoted as $Y_{i,k}\in\mathbb{C}$, for $\forall i\neq k$. The input signal $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is applied by $N$ voltage generators with series admittance $Y_{0}$ to the first $N$ ports of the MiLAC, with $N\leq P$, as represented in Fig. 1. The output signal is read on all $P$ ports, where $P=N+M$ with $M$ denoting the number of ports with no input, and is denoted by $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$, as shown in Fig. 1. For convenience of representation, we introduce the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{p}]^{T}\in\mathbb{C}^{M\times 1}$, such that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$. We consider a MiLAC with $M>0$ ports with no input in the following.
It has been shown in Part I of this paper that the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ can be expressed as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$ by introducing a matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ as
$\mathbf{P}=\frac{\mathbf{Y}}{Y_{0}}+\mathbf{I}_{P},$ (1)
partitioned as
$\mathbf{P}=\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix},$ (2)
where $\mathbf{P}_{11}\in\mathbb{C}^{N\times N}$, $\mathbf{P}_{12}\in\mathbb{C}^{N\times M}$, $\mathbf{P}_{21}\in\mathbb{C}^{M\times N}$, and $\mathbf{P}_{22}\in\mathbb{C}^{M\times M}$. Assuming $\mathbf{P}$ to be invertible, with $\mathbf{P}_{11}$ or $\mathbf{P}_{22}$ invertible, $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are expressed depending on the invertibility of $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ as follows:
  1. 1.
    If $\mathbf{P}_{11}$ is invertible, we have
    $\mathbf{v}_{1}=\Big{(}\mathbf{P}_{11}^{-1}-\mathbf{P}_{11}^{-1}\mathbf{P}_{12}\Big{.}\\ \left.\times\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\right)\mathbf{u},$ (3)
    $\mathbf{v}_{2}=\left(\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{P}_{12}-\mathbf{P}_{22}\right)^{-1}\mathbf{P}_{21}\mathbf{P}_{11}^{-1}\mathbf{u}.$ (4)
  2. 2.
    If $\mathbf{P}_{22}$ is invertible, we have
    $\mathbf{v}_{1}=-\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\mathbf{u},$ (5)
    $\mathbf{v}_{2}=\mathbf{P}_{22}^{-1}\mathbf{P}_{21}\left(\mathbf{P}_{12}\mathbf{P}_{22}^{-1}\mathbf{P}_{21}-\mathbf{P}_{11}\right)^{-1}\mathbf{u}.$ (6)
  3. 3.
    If $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are both invertible, (3)-(4) and (5)-(6) are equivalent.
Note that (3)-(4) and (5)-(6) express the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as a function of the input $\mathbf{u}$ and the blocks of $\mathbf{P}$, which are directly related to $\mathbf{Y}$ by (1). The expressions (3)-(4) and (5)-(6) can be computed by a MiLAC given any blocks $\mathbf{P}_{11}$, $\mathbf{P}_{12}$, $\mathbf{P}_{21}$, and $\mathbf{P}_{22}$ by setting the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ as a function of any arbitrary $\mathbf{P}$ as
$Y_{i,k}=\begin{cases}-Y_{0}\left[\mathbf{P}\right]_{i,k}&i\neq k\\ Y_{0}\sum_{p=1}^{P}\left[\mathbf{P}\right]_{p,k}-Y_{0}&i=k\end{cases},$ (7)
for $i,k=1,\ldots,P$, as discussed in Part I of this paper. Remarkably, a MiLAC can efficiently compute (3)-(4) and (5)-(6) in the analog domain, while they would require computationally expensive matrix-matrix products and matrix inversion operations if digitally computed.
III Analog Computing for the Special Cases of the LMMSE Estimator
In Part I of this paper, we have shown how a MiLAC can be used to compute the LMMSE estimator of a linear observation process with reduced computational complexity. In this section, we derive and discuss five special cases of the LMMSE estimator for linear observation processes, of particular interest in wireless communications. Then, we show how these five special cases can be efficiently computed by a MiLAC.
III-A LMMSE Estimator for Linear Observation Processes
Consider a linear observation process $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{y}\in\mathbb{C}^{Y\times 1}$ is the observation vector, $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. As discussed in the part I of this paper, the LMMSE estimator of $\mathbf{x}$ is equivalently given by
$\hat{\mathbf{x}}_{\textnormal{LMMSE},1}$ $=\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{y},$ (8)
$\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ $=\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{y},$ (9)
where we have $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ when $\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}$ and $\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}$ are both invertible, which holds true in general.
III-B Special Cases of the LMMSE Estimator for Linear Observation Processes
The LMMSE estimator in Section III-A has been obtained with no assumptions on the covariance matrices $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ beyond their invertibility. With specific assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, we can show that it includes five special cases of practical interest, as discussed in the following.
III-B1 Generalized Least Squares
In generalized least squares (GLS), also known as generalized linear regression, the variance on the a priori information on $\mathbf{x}$ is infinitely higher than on $\mathbf{n}$, i.e., $\|\mathbf{C}_{\mathbf{x}}\|_{F}\gg\|\mathbf{C}_{\mathbf{n}}\|_{F}$. Thus, $\mathbf{C}_{\mathbf{x}}^{-1}$ is negligible in (8) and, equivalently, $\mathbf{C}_{\mathbf{n}}$ is negligible in (9), yielding
$\hat{\mathbf{x}}_{\text{GLS},1}$ $=\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{y},$ (10)
$\hat{\mathbf{x}}_{\text{GLS},2}$ $=\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\right)^{-1}\mathbf{y},$ (11)
respectively. Note that, depending on the dimensions of $\mathbf{H}$, only one matrix among $\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}$ and $\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}$ may be invertible. Thus, only one among the estimates $\hat{\mathbf{x}}_{\text{GLS},1}$ and $\hat{\mathbf{x}}_{\text{GLS},2}$ may exist.
III-B2 Generalized Matched Filtering
Opposite to GLS, we can assume that the variance on the a priori information on $\mathbf{x}$ is infinitely smaller than on $\mathbf{n}$, i.e., $\|\mathbf{C}_{\mathbf{x}}\|_{F}\ll\|\mathbf{C}_{\mathbf{n}}\|_{F}$. In this case, that we denote as generalized matched filtering (GMF), both (8) and (9) boil down to
$\hat{\mathbf{x}}_{\text{GMF}}=\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{y}.$ (12)
III-B3 Regularized Least Squares
In regularized least squares (RLS), also known as ridge regression, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, with $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}\in\mathbb{R}$ and $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}>0$. Thus, by introducing $\lambda\in\mathbb{R}$ such that $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}$, (8) and (9) simplify as
$\hat{\mathbf{x}}_{\text{RLS},1}$ $=\left(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{X}\right)^{-1}\mathbf{H}^{H}\mathbf{y},$ (13)
$\hat{\mathbf{x}}_{\text{RLS},2}$ $=\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{Y}\right)^{-1}\mathbf{y},$ (14)
respectively. Interestingly, the expressions in (13) and (14) are widely used in MIMO wireless communications to realize the so-called regularized zero-forcing beamforming (R-ZFBF) and the minimum mean square error (MMSE) receiver, as effective linear transmission and reception techniques, respectively [35, Chapters 7, 12].
TABLE I: Special cases of the LMMSE estimator and their assumptions.
Assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$
LMMSE None
GLS $\|\mathbf{C}_{\mathbf{x}}\|_{F}\gg\|\mathbf{C}_{\mathbf{n}}\|_{F}$
GMF $\|\mathbf{C}_{\mathbf{x}}\|_{F}\ll\|\mathbf{C}_{\mathbf{n}}\|_{F}$
RLS $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$
OLS $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, $\lambda_{\mathbf{x}}\gg\lambda_{\mathbf{n}}$
OMF $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, $\lambda_{\mathbf{x}}\ll\lambda_{\mathbf{n}}$
III-B4 Ordinary Least Squares
In ordinary least squares (OLS), also simply known as LS, or linear regression, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, and the variance on the a priori information on $\mathbf{x}$ is much higher than on $\mathbf{n}$, i.e., $\lambda_{\mathbf{x}}\gg\lambda_{\mathbf{n}}$. Thus, since $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}\ll 1$, (13) and (14) further simplify as
$\hat{\mathbf{x}}_{\text{OLS},1}$ $=\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{y},$ (15)
$\hat{\mathbf{x}}_{\text{OLS},2}$ $=\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\mathbf{y},$ (16)
respectively. Interestingly, OLS is a special case of both GLS and RLS as it includes both their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. As discussed for the GLS, since only one matrix among $\mathbf{H}^{H}\mathbf{H}$ and $\mathbf{H}\mathbf{H}^{H}$ may be invertible depending on the dimensions of $\mathbf{H}$, only one among $\hat{\mathbf{x}}_{\text{OLS},1}$ and $\hat{\mathbf{x}}_{\text{OLS},2}$ may exist, unless $\mathbf{H}$ is a square matrix. Remarkably, the expressions of the OLS are widely employed in MIMO wireless communications to implement zero-forcing beamforming (ZFBF) and the ZF receiver [35, Chapters 7, 12].
III-B5 Ordinary Matched Filtering
In ordinary matched filtering (OMF), also simply known as MF, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, and the variance on the a priori information on $\mathbf{x}$ is much smaller than on $\mathbf{n}$, i.e., $\lambda_{\mathbf{x}}\ll\lambda_{\mathbf{n}}$. Thus, since $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}\gg 1$, both (13) and (14) boil down to
$\hat{\mathbf{x}}_{\text{OMF}}=\lambda^{-1}\mathbf{H}^{H}\mathbf{y}.$ (17)
Remarkably, OMF is a special case of both GMF and RLS as it jointly accounts for their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. In addition, it is utilized in MIMO wireless communications to realize matched beamforming (MBF) and the MF receiver [35, Chapters 7, 12].
We summarize the five special cases of the LMMSE estimator for linear observation processes and their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ in Tab. I.
TABLE II: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{11}$ invertible.
$\mathbf{v}_{2}$ $\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix}$
LMMSE $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GLS $\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\left(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{X}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$
OLS $\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mathbf{0}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\lambda\mathbf{I}_{X}\end{bmatrix}$
III-C Computing the Special Cases of the LMMSE Estimator with a MiLAC
The expressions of the five special cases of the LMMSE estimator can be efficiently returned on the output vector $\mathbf{v}_{2}$ of a MiLAC having $N=Y$ input ports and $M=X$ output ports with no input, i.e., $P=X+Y$ total ports. This is achieved by exploiting (4) and (6), as discussed for the LMMSE estimator in Part I of this paper. The input vector of the MiLAC is set as $\mathbf{u}=\mathbf{y}$ and the tunable admittance components $\{Y_{i,k}\}_{i,k=1}^{P}$ are set according to (7) such that the blocks of the matrix $\mathbf{P}$ give the desired expression in $\mathbf{v}_{2}$. In detail, the expressions of the special cases in (10)-(17) can be computed by setting the matrix $\mathbf{P}$ as illustrated in Tabs. II and III, depending on the invertibility of the blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$. When the block $\mathbf{P}_{11}$ is invertible, $\mathbf{v}_{2}$ is given by (4), as reported in Tab. II, while when the block $\mathbf{P}_{22}$ is invertible, $\mathbf{v}_{2}$ is given by (6), as reported in Tab. III. Note that GLS and OLS are computed via a matrix $\mathbf{P}$ whose blocks $\mathbf{P}_{11}$ and $\mathbf{P}_{22}$ are not both invertible. Thus, for GLS and OLS the expressions in Tabs. II and III are not equivalent in general, while they are equivalent for the LMMSE estimator and the other special cases.
Remark 1.
The five special cases of the LMMSE estimator can be efficiently computed using a MiLAC, which offers significantly reduced computational complexity compared to a digital computer. This is because a MiLAC directly performs matrix-matrix multiplications and matrix inversion operations in the analog domain. On the one hand, the computational complexity of computing the special cases with a MiLAC is driven by the complexity of computing (7) for a given matrix $\mathbf{P}$. Thus, the complexity required to compute GLS, RLS, and OLS with a MiLAC is the same as the complexity of computing the LMMSE estimator, i.e., $6XY$ real operations, as discussed in Part I of this paper. Besides, computing (7) for the GMF and OMF require approximately $2XY$ real operations for both cases $i\neq k$ and $i=k$, resulting in a total of $4XY$ real operations, where we assumed that the operations involving $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are precomputed offline. On the other hand, a digital computer calculates the GLS with the same complexity as the LMMSE estimator due to the two matrix-matrix products and matrix inversion, i.e., $8(XY^{2}+X^{2}Y+\min\{X^{3},Y^{3}\}/3)$, as discussed in Part I of this paper. The RLS and OLS can be computed with reduced complexity as they require only one matrix-matrix product and a matrix inversion, namely $8\min\{X^{2}Y+X^{3}/3,XY^{2}+Y^{3}/3\}$. Besides, the GMF and OMF only require three and one matrix-vector products, respectively, resulting in a complexity of $8(X^{2}+XY+Y^{2})$ and $8XY$ real operations, respectively.
TABLE III: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{22}$ invertible.
$\mathbf{v}_{2}$ $\begin{bmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}\\ \mathbf{P}_{21}&\mathbf{P}_{22}\end{bmatrix}$
LMMSE $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}\right)^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GLS $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\mathbf{0}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
GMF $\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\mathbf{C}_{\mathbf{n}}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{C}_{\mathbf{x}}^{-1}\end{bmatrix}$
RLS $\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{Y}\right)^{-1}\mathbf{u}$ [±$\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
OLS $\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\mathbf{u}$ $\begin{bmatrix}\mathbf{0}_{Y}&\mathbf{H}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
OMF $\lambda^{-1}\mathbf{H}^{H}\mathbf{u}$ [±$\begin{bmatrix}\pm\lambda\mathbf{I}_{Y}&\mathbf{0}_{Y\times X}\\ \mathbf{H}^{H}&\mp\mathbf{I}_{X}\end{bmatrix}$
IV A Novel Beamforming Strategy:
MiLAC-Aided Beamforming
In this section, we show how a MiLAC can be used to implement a novel beamforming technique, that we denote as MiLAC-aided beamforming, at the transmitter as well as receiver side. MiLAC-aided beamforming has the same flexibility as digital beamforming, while it offers circuit and computational complexity benefits by fully operating in the analog domain. To illustrate MiLAC-aided beamforming, we consider a MIMO communication system between $N_{T}$ transmitting antennas and $N_{R}$ receiving antennas through the wireless channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$. We denote as $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ the transmitted signal and as $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ the received signal such that $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{n}\in\mathbb{C}^{N_{R}\times 1}$ is the additive white Gaussian noise (AWGN)  [^1] We consider here a general MIMO system, such as single-user MIMO or multi-user with one symbol per user. Both single- and multi-user systems will be specifically studied in the following subsections.. We begin by showing how MiLAC-aided beamforming can be used to implement an arbitrary beamforming matrix.
Figure 2: MiLAC-aided beamforming at the transmitter, performing $\mathbf{x}=\mathbf{W}\mathbf{s}$ in the analog domain, where $\mathbf{s}$ is the symbol vector and $\mathbf{W}$ is an arbitrary precoding matrix.
IV-A Arbitrary Beamforming
Consider a transmitter that linearly precodes a vector of $N_{S}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$ to obtain the transmitted signal as $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{S}}$ is an arbitrary precoding matrix. Such a transmitter can be implemented purely in the analog domain by a MiLAC having $N=N_{S}$ ports receiving the symbol vector in input and $M=N_{T}$ ports delivering the output to antennas perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, as represented in Fig. 2. Since the antennas are perfectly matched, their Thevenin equivalent circuit is an impedance $Z_{0}=Y_{0}^{-1}$ connected to ground [36, Chapter 2.13]. Thus, the MiLAC in Fig. 2 behaves as when its last $N_{T}$ ports are connected to ground through an impedance $Z_{0}=Y_{0}^{-1}$, i.e., as analyzed in Section II. Following the analysis in Section II, such a MiLAC returns $\mathbf{x}=\mathbf{W}\mathbf{s}$ on its last $N_{T}$ ports when the signal $\mathbf{s}$ is fed as input on the first $N_{S}$ ports and the tunable admittance components are set such that the matrix $\mathbf{P}$ is
$\mathbf{P}=\begin{bmatrix}\mathbf{I}_{N_{S}}&\mathbf{0}_{N_{S}\times N_{T}}\\ -\mathbf{W}&\mathbf{I}_{N_{T}}\end{bmatrix},$ (18)
according to (4) and (6). To this end, the tunable admittance components are set based on (7) where $\mathbf{P}$ is given by (18), yielding
$Y_{i,k}=\begin{cases}Y_{0}\left[\mathbf{W}\right]_{i-N,k}&N_{S}<i,\;k\leq N_{S}\\ 0&\text{Otherwise}\end{cases},$ (19)
for $i\neq k$, and
$Y_{k,k}=\begin{cases}-Y_{0}\sum_{n_{T}=1}^{N_{T}}\left[\mathbf{W}\right]_{n_{T},k}&k\leq N_{S}\\ 0&k>N_{S}\end{cases},$ (20)
for $k=1,\ldots,N_{S}+N_{T}$. From (19) and (20), we observe that the tunable admittance components of a MiLAC can be reconfigured in closed form to implement any arbitrary beamforming $\mathbf{W}$ matrix in the analog domain. The computational complexity of this beamforming technique, denoted as MiLAC-aided beamforming, is given by the complexity of designing $\mathbf{W}$ (depending on the adopted precoding strategy) and computing (19) and (20) (requiring in total $4N_{S}N_{T}$ real operations). Note that these operations are executed at each coherence block, while no operation is required on a per-symbol basis since the matrix-vector product $\mathbf{W}\mathbf{s}$ is performed in the analog domain.
Figure 3: MiLAC-aided beamforming at the receiver, performing $\mathbf{z}=\mathbf{G}\mathbf{y}$ in the analog domain, where $\mathbf{y}$ is the received signal and $\mathbf{G}$ is an arbitrary combining matrix.
As discussed on the transmitter side, MiLAC-aided beamforming can also be introduced on the receiver side. Specifically, a MiLAC can be used to linearly combine the received signal $\mathbf{y}\in\mathbb{C}^{N_{R}\times 1}$ to obtain the signal used for detection $\mathbf{z}\in\mathbb{C}^{N_{S}\times 1}$ as $\mathbf{z}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}\in\mathbb{C}^{N_{S}\times N_{R}}$ is an arbitrary combining matrix. This can be obtained with MiLAC having $N=N_{R}$ ports acquiring the input from the receiving antennas and $M=N_{S}$ ports providing the output vector $\mathbf{z}$, as represented in Fig. 3. Note that in Fig. 3 we assume the antennas to be perfectly matched to the reference impedance $Z_{0}=Y_{0}^{-1}$, such that their Thevenin equivalent circuit is a voltage generator (inducing a voltage $y_{n}$ at port $n$, for $n=1,\ldots,N$) in series to an impedance $Z_{0}=Y_{0}^{-1}$ [36, Chapter 2.13], i.e., as in the MiLAC analyzed in Section II.
Among the possible linear precoding and combining strategies, MiLAC-aided beamforming can efficiently perform R-ZFBF, ZFBF, and MBF at the transmitter and MMSE, ZF, and MF at the receiver by exploiting the capability of a MiLAC to compute the corresponding special case of the LMMSE estimator. This leads to remarkable computational complexity benefits, as illustrated in the following.
IV-B LMMSE-Inspired Beamforming at the Transmitter:
R-ZFBF, ZFBF, and MBF Transmitters
Assume that the previously considered MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a multi-user system between an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna users, with $N_{R}\leq N_{T}$. The signal received at the $N_{R}$ users $\mathbf{y}$ is used to detect the symbols $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$, with $N_{S}=N_{R}$, and the transmitted signal is given by $\mathbf{x}=\mathbf{W}\mathbf{s}$, where $\mathbf{W}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the precoding matrix implemented by a MiLAC, as in Fig. 2. By exploiting the operations included in Tab. III, it is possible to reconfigure the microwave network of the MiLAC to efficiently execute three popular linear transmitters [35, Chapter 12]. First, a MiLAC computing RLS can realize R-ZFBF, with precoding matrix $\mathbf{W}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$, where $\lambda$ is an arbitrary regularization factor. Second, ZFBF can be implemented by a MiLAC computing OLS, whose precoding matrix is $\mathbf{W}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$. Third, a MiLAC computing OMF can realize MBF, with precoding matrix given by $\mathbf{W}_{\text{MBF}}=\lambda^{-1}\mathbf{H}^{H}$. These three transmitters are obtained by reconfiguring the tunable admittance components of the MiLAC according to the channel $\mathbf{H}$ on a per coherence block basis, and do not require any operation on a per-symbol basis.
Remark 2.
According to Remark 1, the R-ZFBF and ZFBF transmitters can be achieved with only $6N_{T}N_{R}$ real operations, while the MBF transmitter with only $4N_{T}N_{R}$ real operations, which are necessary to set the tunable admittance components of the MiLAC. Conversely, the computational complexity of digitally calculating the R-ZFBF and ZFBF transmitters is $8(N_{T}N_{R}^{2}+N_{R}^{3}/3)$ real operations, while for the MBF transmitter is $8N_{T}N_{R}$. This confirms the significant benefits of MiLAC-aided beamforming in terms of computational complexity, especially in implementing the R-ZFBF and ZFBF transmitters.
IV-C LMMSE-Inspired Beamforming at the Receiver:
MMSE, ZF, and MF Receivers
Assume that the MIMO system $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$ is a single-user system between an $N_{T}$-antenna transmitter and a $N_{R}$-antenna receiver, with $N_{R}\geq N_{T}$. The transmitted signal $\mathbf{x}$ contains the symbols to be detected at the receiver, i.e., $\mathbf{x}=\mathbf{s}$, and the receiver uses the signal $\mathbf{z}=\mathbf{G}\mathbf{y}\in\mathbb{C}^{N_{T}\times 1}$ to accurately detect the symbols, where $\mathbf{G}\in\mathbb{C}^{N_{T}\times N_{R}}$ is the combining matrix imposed by the MiLAC, as in Fig. 3. By using the operations in Tab. II, it is possible to reconfigure the microwave network of the MiLAC to efficiently implement three widely used receivers [35, Chapter 7]. First, a MiLAC computing RLS can realize the MMSE receiver with combining matrix $\mathbf{G}_{\text{MMSE}}=(\mathbf{H}^{H}\mathbf{H}+\lambda\mathbf{I}_{N_{T}})^{-1}\mathbf{H}^{H}$. If the covariance matrix of the transmitted signal $\mathbf{x}$ is $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$, where $P_{T}$ is the transmit power, and the covariance matrix of the noise $\mathbf{n}$ is $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$, where $\sigma_{\mathbf{n}}^{2}$ is the noise power, the optimal regularization factor $\lambda$ is given by $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ [35, Chapter 7]. Second, the ZF receiver can be realized by a MiLAC computing OLS, whose combining matrix is $\mathbf{G}_{\text{ZF}}=(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{H}^{H}$. Third, a MiLAC computing OMF can implement the MF receiver, with combining matrix $\mathbf{G}_{\text{MF}}=\lambda^{-1}\mathbf{H}^{H}$, where the optimal $\lambda$ is $\lambda=N_{T}\sigma_{\mathbf{n}}^{2}/P_{T}$ if $\mathbf{C}_{\mathbf{x}}=P_{T}/N_{T}\mathbf{I}_{N_{T}}$ and $\mathbf{C}_{\mathbf{n}}=\sigma_{\mathbf{n}}^{2}\mathbf{I}_{N_{R}}$ [35, Chapter 7]. These three receivers are realized by solely reconfiguring the MiLAC on a per coherence block basis, while no operation is required on a per-symbol basis.
Remark 3.
Similarly to Remark 2, the MMSE and ZF receivers can be obtained through a MiLAC with only $6N_{T}N_{R}$ real operations and the MF receiver with only $4N_{T}N_{R}$ real operations, which are required to set the tunable admittance components. In contrast, the computational complexity of digitally calculating the MMSE and ZF receivers is $8(N_{T}^{2}N_{R}+N_{T}^{3}/3)$ real operations, and for the MF receiver is $8N_{T}N_{R}$, showing important benefits in using MiLAC-aided beamforming to implement the MMSE and ZF receivers.
IV-D Computing the DFT
MiLAC-aided beamforming is also beneficial when used to precode or combine a signal with a fixed beamforming matrix. In this case, since the admittance components of the MiLAC are optimized offline, no operation is required online with MiLAC-aided beamforming, while a matrix-vector product is needed on a per-symbol basis in the case of digital beamforming. Combining the received signal $\mathbf{y}$ with a fixed matrix $\mathbf{G}$ is particularly useful for DoA estimation, where the received signal is typically multiplied by the DFT matrix [30]. This operation performed at the receiver can be formalized as $\mathbf{z}=\mathbf{G}\mathbf{y}$, where $\mathbf{G}\in\mathbb{C}^{N_{R}\times N_{R}}$ is the DFT matrix, given by
$\left[\mathbf{G}\right]_{i,k}=\frac{1}{\sqrt{N_{R}}}e^{-j2\pi\frac{(i-1)(k-1)}{N_{R}}},$ (21)
for $i,k=1,\ldots,N_{R}$. This operation can be implemented purely in the analog domain by a MiLAC having $N=N_{R}$ ports acquiring the input $\mathbf{y}$ from the receiving antennas and $M=N_{R}$ ports returning the output, as represented in Fig. 3. To perform such a DFT operation, the admittance components of the MiLAC need to be set offline following (7) where the matrix $\mathbf{P}$ given by
$\mathbf{P}=\begin{bmatrix}\mathbf{I}_{N_{R}}&\mathbf{0}_{N_{R}\times N_{R}}\\ -\mathbf{G}&\mathbf{I}_{N_{R}}\end{bmatrix},$ (22)
with $\mathbf{G}$ being the DFT matrix in (21). Thus, by substituting (21) and (22) into (7), we obtain that the admittance components $\{Y_{i,k}\}_{i,k=1}^{2N_{R}}$ are set as
$Y_{i,k}=\begin{cases}\frac{Y_{0}}{\sqrt{N_{R}}}e^{-j2\pi\frac{(i-N-1)(k-1)}{N_{R}}}&N_{R}<i,\;k\leq N_{R}\\ 0&\text{Otherwise}\end{cases},$ (23)
for $i\neq k$, and
$Y_{k,k}=\begin{cases}-Y_{0}\sqrt{N_{R}}&k=1\\ 0&k\neq 1\end{cases},$ (24)
for $k=1,\ldots,2N_{R}$.
Remark 4.
By employing a MiLAC with fixed admittance components set as in (23) and (24), we can perform the DFT instantly in the analog domain requiring no digital operations. Conversely, the DFT requires approximately $34/9N_{R}\log_{2}(N_{R})$ real operations to be digitally computed [37], which could become prohibitive as the number of antennas $N_{R}$ grows extremely large. Note that we assume a one-dimensional DFT for simplicity, while a similar discussion also holds to perform a two-dimensional DFT in the analog domain, as considered in [30].
Figure 4: (a) Digital, (b) analog, (c) hybrid, (d) RIS-aided, (e) SIM-aided, and (f) MiLAC-aided beamforming at the transmitter.
TABLE IV: Beamforming strategies comparison.
Beamforming Precoding matrix Constraints
Number of
RF chains
Resolution of
ADCs/DACs
Operations per
symbol basis
Operations per
coherence block
Digital $\mathbf{W}_{D}$ $\mathbf{W}_{D}\in\mathbb{C}^{N_{T}\times N_{S}}$ $N_{RF}=N_{T}$ High Compute $\mathbf{W}_{D}\mathbf{s}$ Optimize $\mathbf{W}_{D}$
Analog $\mathbf{w}_{A}$ $\mathbf{w}_{A}\in\mathbb{T}^{N_{T}\times 1}$ $N_{RF}=N_{S}=1$ Low None Optimize $\mathbf{w}_{A}$
Hybrid $\mathbf{W}_{A}\mathbf{W}_{D}$
$\mathbf{W}_{D}\in\mathbb{C}^{N_{RF}\times N_{S}}$
$\mathbf{W}_{A}\in\mathbb{T}^{N_{T}\times N_{RF}}$
$N_{RF}\geq N_{S}$ High Compute $\mathbf{W}_{D}\mathbf{s}$ Optimize $\mathbf{W}_{D},\mathbf{W}_{A}$
RIS-aided $\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}\mathbf{W}_{D}$
$\mathbf{W}_{D}\in\mathbb{C}^{N_{RF}\times N_{S}}$
$\mathbf{H}_{\text{RIS}}\in\mathbb{C}^{N_{T}\times N_{RF}}$, fixed
$\boldsymbol{\Theta}\in\mathbb{C}^{N_{T}\times N_{T}},\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}_{N_{T}}$
$N_{RF}\geq N_{S}$ High Compute $\mathbf{W}_{D}\mathbf{s}$ Optimize $\mathbf{W}_{D},\boldsymbol{\Theta}$
SIM-aided
$\boldsymbol{\Theta}_{L}\mathbf{H}_{\text{SIM},L}\cdots$
$\cdots\boldsymbol{\Theta}_{1}\mathbf{H}_{\text{SIM},1}$
$\mathbf{H}_{\text{SIM},1}\in\mathbb{C}^{N_{T}\times N_{RF}}$, fixed
$\mathbf{H}_{\text{SIM},\ell}\in\mathbb{C}^{N_{T}\times N_{T}}$, $\ell>2$, fixed
$\boldsymbol{\Theta}_{\ell}\in\mathbb{C}^{N_{T}\times N_{T}},\boldsymbol{\Theta}_{\ell}^{H}\boldsymbol{\Theta}_{\ell}=\mathbf{I}_{N_{T}}$
$N_{RF}=N_{S}$ Low None Optimize $\boldsymbol{\Theta}_{\ell}$
MiLAC-aided $\mathbf{W}_{A}$ $\mathbf{W}_{A}\in\mathbb{C}^{N_{T}\times N_{S}}$ $N_{RF}=N_{S}$ Low None Optimize $Y_{i,k}$
V Comparison with Existing Beamforming Strategies
We have discussed how a MiLAC can enable MiLAC-aided beamforming, a new beamforming strategy purely in the analog domain having the same flexibility as digital beamforming. In this section, we highlight the benefits of MiLAC-aided beamforming over the state-of-the-art beamforming strategies, showing that it requires fewer RF chains, lower-resolution ADCs/DACs, and reduced computational complexity. To this end, we consider a transmitter with $N_{T}$ antennas and $N_{RF}$ RF chains transmitting $N_{S}$ symbols, where we have $N_{S}\leq N_{RF}\leq N_{T}$ [^2] Several strategies have been proposed to perform symbol modulation in the analog domain, i.e., through ESPAR [22, 23, 24], LMA [25, 26], and RIS [27, 28], which can enable multi-stream transmissions ($N_{S}>1$) with $N_{RF}=1$ RF chain. However, these strategies operate symbol modulation on a per-symbol basis, and are fundamentally different from beamforming strategies acting per coherence block. . We denote as $\mathbf{s}\in\mathbb{C}^{N_{S}\times 1}$ the symbol vector and as $\mathbf{x}\in\mathbb{C}^{N_{T}\times 1}$ the transmitted signal. The following discussion for a transmitter readily applies also for a receiver.
V-A Digital Beamforming
In digital beamforming, the transmitted signal is given by $\mathbf{x}=\mathbf{W}_{D}\mathbf{s}$, where $\mathbf{W}_{D}\in\mathbb{C}^{N_{T}\times N_{S}}$ is the digital precoder, subject only to the power constraint. To allow maximum flexibility, there are $N_{RF}=N_{T}$ RF chains equipped high-resolution ADCs/DACs, each carrying an entry of the transmitted signal $\mathbf{x}$ to the corresponding antenna, as in Fig. 4(a). In addition, the computational complexity is driven by the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ executed on a per-symbol basis and by the optimization of $\mathbf{W}_{D}$ at each coherence block based on the channel realization, whose complexity depends on the specific digital precoding strategy.
V-B Analog Beamforming
In analog beamforming [5], only one symbol $s\in\mathbb{C}$ is transmitted ($N_{S}=1$), and the transmitted signal is given by $\mathbf{x}=\mathbf{w}_{A}s$, where $\mathbf{w}_{A}\in\mathbb{C}^{N_{T}\times 1}$ is the analog precoder, as in Fig. 4(b). Since the symbol $s$ reaches the $N_{T}$ antennas by passing through $N_{T}$ phase shifters, the analog precoder is subject to $\mathbf{w}_{A}\in\mathbb{T}^{N_{T}\times 1}$, where $\mathbb{T}=\{c\in\mathbb{C}:|c|=1\}$, due to the unit-modulus constraint imposed by the phase shifters. In this case, a single RF chain is sufficient to carry the symbol $s$, which can be equipped with low-resolution ADCs/DACs since the symbol $s$ lies in a constellation with finite cardinality. The computational complexity of analog beamforming is given by the complexity of optimizing $\mathbf{w}_{A}$ at every coherence block as a function of the channel realization. Note that no operation takes place on a per-symbol basis since the product $\mathbf{w}_{A}s$ is performed in the analog domain as the signal flows through the phase shifters connected to all antennas.
V-C Hybrid Beamforming
In hybrid beamforming [5, 6, 7, 8, 9], the transmitted signal is $\mathbf{x}=\mathbf{W}_{A}\mathbf{W}_{D}\mathbf{s}$, where $\mathbf{W}_{A}\in\mathbb{C}^{N_{T}\times N_{RF}}$ is the analog precoder, subject to specific constraints depending on the architecture of phase shifters and switches, and $\mathbf{W}_{D}\in\mathbb{C}^{N_{RF}\times N_{S}}$ is the digital precoder, subject only to the power constraint, as in Fig. 4(c). Thus, the effective precoding matrix is given by $\mathbf{W}=\mathbf{W}_{A}\mathbf{W}_{D}$. Numerous hybrid beamforming architectures have been proposed, each with different constraints on the matrix $\mathbf{W}_{A}$. For example, in the fully-connected architecture, each RF chain is connected to all the antennas through a phase shifter, and each entry of $\mathbf{W}_{A}$ is subject to the unit-modulus constraint, i.e., $\mathbf{W}_{A}\in\mathbb{T}^{N_{T}\times N_{RF}}$. The computational complexity of hybrid beamforming is given by the complexity of computing the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ on a per-symbol basis, and optimizing $\mathbf{W}_{A}$ and $\mathbf{W}_{D}$ at each coherence block.
V-D RIS-Aided Beamforming
More recently, RIS-aided beamforming emerged by exploiting a RIS deployed in close proximity with an active transmitting device [12, 13, 14, 18], as in Fig. 4(d). The RIS can work in either reflective or transmissive mode, without changing the system model. Denoting as $\mathbf{x}^{\prime}=\mathbf{W}_{D}\mathbf{s}$ the digitally precoded signal transmitted by the active device, the received signal is given by $\mathbf{y}=\mathbf{H}_{\text{EFF}}\mathbf{x}^{\prime}$, where $\mathbf{H}_{\text{EFF}}\in\mathbb{C}^{N_{R}\times N_{RF}}$ is the RIS-aided channel, given by $\mathbf{H}_{\text{EFF}}=\mathbf{H}\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}$, with $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$, $\boldsymbol{\Theta}\in\mathbb{C}^{N_{T}\times N_{T}}$, and $\mathbf{H}_{\text{RIS}}\in\mathbb{C}^{N_{T}\times N_{RF}}$ being the channel from the RIS to the receiver, the scattering matrix of the RIS, and the channel from the active transmitting device to the RIS, respectively. Thus, effective precoding matrix is $\mathbf{W}=\boldsymbol{\Theta}\mathbf{H}_{\text{RIS}}\mathbf{W}_{D}$, which is optimized by reconfiguring $\boldsymbol{\Theta}$, subject to $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}_{N_{T}}$ in the case of a lossless RIS [15], and $\mathbf{W}_{D}$, subject only to the power constraint. As in hybrid beamforming, the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, while $\mathbf{W}_{D}$ and $\boldsymbol{\Theta}$ are optimized per coherence block.
V-E SIM-Aided Beamforming
RIS-aided beamforming has been extended by stacking multiple transmissive RISs at the transmitter, which is denoted in the literature as a SIM [19, 20, 21], as in Fig. 4(e). Given the additional flexibility provided by the multiple metasurfaces, SIM-aided beamforming does not include any digital precoding, and for this reason it only requires $N_{RF}=N_{S}$ RF chains equipped with low-resolution ADCs/DACs [19, 20, 21, 30]. Similar to RIS-aided beamforming, SIM-aided beamforming exploits the fact that the effective channel between the transmitter and receiver can be engineered by reconfiguring the scattering matrices of the stacked RISs. Assuming a SIM made of $L$ stacked RISs each with $N_{T}$ elements, the effective channel between the active device transmitting the symbols $\mathbf{s}$ and the receiver is given by
$\mathbf{H}_{\text{EFF}}=\mathbf{H}\boldsymbol{\Theta}_{L}\mathbf{H}_{\text{SIM},L}\cdots\boldsymbol{\Theta}_{2}\mathbf{H}_{\text{SIM},2}\boldsymbol{\Theta}_{1}\mathbf{H}_{\text{SIM},1},$ (25)
where $\mathbf{H}\in\mathbb{C}^{N_{R}\times N_{T}}$, $\boldsymbol{\Theta}_{\ell}\in\mathbb{C}^{N_{T}\times N_{T}}$, and $\mathbf{H}_{\text{SIM},\ell}\in\mathbb{C}^{N_{T}\times N_{T}}$ are the channel from the $L$th RIS to the receiver, the scattering matrix of the $\ell$th RIS, and the channel from the $\ell-1$th to the $\ell$th RIS (where the 00th RIS is the transmitting device and $\mathbf{H}_{\text{SIM},1}\in\mathbb{C}^{N_{T}\times N_{RF}}$), respectively. Thus, the effective precoding matrix reads as
$\mathbf{W}=\boldsymbol{\Theta}_{L}\mathbf{H}_{\text{SIM},L}\cdots\boldsymbol{\Theta}_{2}\mathbf{H}_{\text{SIM},2}\boldsymbol{\Theta}_{1}\mathbf{H}_{\text{SIM},1},$ (26)
which is optimized per coherence block by reconfiguring the scattering matrices of the $L$ stacked RISs $\boldsymbol{\Theta}_{\ell}$, for $\ell=1,\ldots,L$.
Since SIM-aided beamforming is entirely performed in the analog domain, it can be regarded as a low-complexity approach for implementing MiLAC-aided beamforming, with fewer tunable admittance components. Due to the reduced number of tunable components, the flexibility of the beamforming matrix is consequently reduced, and it is constrained as in (26) and cannot be an arbitrary matrix.
V-F MiLAC-Aided Beamforming
In MiLAC-aided beamforming, the symbols $\mathbf{s}$ are given in input to a MiLAC, and the transmitted signal $\mathbf{x}$ is returned on its output ports, as explained in Section IV. This novel beamforming strategy brings five unique benefits over state-of-the-art beamforming strategies. Specifically, it offers the maximum flexibility with minimum hardware complexity (in terms of the number of RF chains and resolution of ADCs/DACs) and minimum computational complexity (in terms of operations per-symbol basis and per coherence block).
V-F1 Maximum Flexibility
MiLAC-aided beamforming has the same flexibility as digital beamforming since it can apply any arbitrary beamforming matrix $\mathbf{W}$, as illustrated in Section IV. Given its full flexibility, MiLAC-aided beamforming can obtain the same performance as digital beamforming, outperforming other beamforming strategies that impose limiting constraints on $\mathbf{W}$, i.e., analog, hybrid, RIS-aided, and SIM-aided, as summarized in Tab. IV.
V-F2 Minimum Number of RF Chains
In MiLAC-aided beamforming, the symbols are fed in input to the MiLAC, which computes the transmitted signal in the analog domain. Thus, only $N_{RF}=N_{S}$ RF chains are needed to achieve the full flexibility of digital beamforming, which is the minimum number of RF chains requested by the reviewed beamforming strategies, as shown in Tab. IV. Note that a fully-connected hybrid beamforming architecture requires $N_{RF}=2N_{S}$ RF chains to realize any digital beamforming matrix [7], while the other beamforming strategies are not proven to achieve full flexibility with any $N_{RF}<N_{T}$.
V-F3 Low-resolution ADCs/DACs
Since MiLAC-aided beamforming processes the symbols purely in the analog domain, each RF chain carries an individual symbol, which lies in a constellation with finite cardinality. Thus, low-resolution ADCs/DACs can be adopted without sacrificing performance, which are less expensive and less power-hungry than high-resolution ADCs/DACs. High-resolution ADCs/DACs are instead necessary when beamforming is performed in the digital domain (entirely or in part), i.e., in digital beamforming, hybrid beamforming, and RIS-aided beamforming, as indicated in Tab. IV.
V-F4 No Computations on a Per-Symbol Basis
If beamforming is performed in the digital domain (entirely or in part), the matrix-vector product $\mathbf{W}_{D}\mathbf{s}$ needs to be computed on a per-symbol basis, as shown in Tab. IV. The complexity of this operation is $8N_{RF}N_{S}$ real operations, as discussed in the Appendix of Part I of this paper, and can become prohibitive in massive/gigantic MIMO systems [3]. MiLAC-aided beamforming solves this issue since it does not require any computation on a per-symbol basis, as the beamforming is purely analog. This reduces the beamforming computational complexity, especially when the coherence time is much longer than the symbol time [3].
V-F5 Minimum Computations Per Coherence Block for RLS and OLS
MiLAC-aided beamforming offers further gains in terms of computational complexity by exploiting the capability of a MiLAC to efficiently compute RLS and OLS in the analog domain. Specifically, these gains can be obtained when MiLAC-aided beamforming is used to compute the R-ZFBF and ZFBF transmitters or the MMSE and ZF receivers, as discussed in Section IV.
Figure 5: Sum rate in a $4\times 4$ multi-user MISO system, with digital and MiLAC-aided beamforming at the transmitter.
VI Numerical Results
We have proposed MiLAC-aided beamforming as a novel beamforming strategy that offers the same flexibility as digital beamforming, while presenting promising advantages in terms of circuit and computational complexity. In this section, we evaluate the performance and computational complexity of MiLAC-aided beamforming.
VI-A Performance Evaluation
Since MiLAC-aided beamforming can implement any arbitrary beamforming matrix, it is proven to perform as digital beamforming and outperform other beamforming strategies that are limited by constrained beamforming matrices. The specific performance gains over the reviewed beamforming strategies match those of digital beamforming, which depend on the performance metric and scenario, such as whether the transmission is single- or multi-stream, and whether the system is single- or multi-user. Thus, we analyze the cases where MiLAC-aided beamforming leverages the computing capabilities of a MiLAC to implement the LMMSE-inspired beamforming matrices, i.e., R-ZFBF, ZFBF, and MBF at the transmitter, and the MMSE, ZF, and MF at the receiver.
Figure 6: BER in a $4\times 4$ single-user MIMO system using QPSK, with digital and MiLAC-aided beamforming at the receiver.
To evaluate the performance of R-ZFBF, ZFBF, and MBF realized with digital and MiLAC-aided beamforming, we consider a multi-user multiple-input single-output (MISO) system with an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna receivers, as presented in Section IV-B. Here, the transmitter precodes the $N_{R}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{R}\times 1}$ intended to the $N_{R}$ receivers through R-ZFBF or ZFBF to suppress inter-stream interference, or through MBF to maximize the channel gains. With digital beamforming, the R-ZFBF transmitter is realized by computing $\mathbf{F}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$ and setting the precoding matrix $\mathbf{W}_{\text{R-ZFBF}}$ such that $[\mathbf{W}_{\text{R-ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, assuming uniform power allocation to the users. For large $N_{R}$ and if the signal-to-noise ratio (SNR) is the same at all receiving antennas, the optimal $\lambda$ is $\lambda=N_{R}\sigma_{\mathbf{n}}^{2}/P_{T}$, where $P_{T}$ is the transmit power, i.e., $\mathbf{s}$ has covariance matrix $\mathbf{C}_{\mathbf{s}}=P_{T}/N_{R}\mathbf{I}_{N_{R}}$, $\sigma_{\mathbf{n}}^{2}$ is the noise power, and we assumed the channel to have unit path gain [35, Chapter 12]. Similarly, digital beamforming can realize ZFBF by computing $\mathbf{F}_{\text{ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ and setting $\mathbf{W}_{\text{ZFBF}}$ as $[\mathbf{W}_{\text{ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, and MBF by using $\mathbf{F}_{\text{MBF}}=\mathbf{H}^{H}$ and $\mathbf{W}_{\text{MBF}}$ such that $[\mathbf{W}_{\text{MBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{MBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$. When a MiLAC is used to perform R-ZFBF, ZFBF, and MBF, we have two choices. First, the MiLAC can be reconfigured to implement an arbitrary beamforming matrix, as discussed in Section IV-A. Thus, its tunable admittance components can be set according to (19) and (20), where $\mathbf{W}$ is the desired beamforming matrix $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, or $\mathbf{W}_{\text{MBF}}$ with uniform power allocation, which is digitally computed. Second, the MiLAC can directly compute the precoding matrices of these LMMSE-inspired beamforming techniques in the analog domain, as explained in Section IV-B. In this case, the columns of $\mathbf{W}_{\text{R-ZFBF}}$, $\mathbf{W}_{\text{ZFBF}}$, and $\mathbf{W}_{\text{MBF}}$ cannot be individually normalized, and the resulting precoding matrices are given by $\mathbf{W}_{\text{R-ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{R-ZFBF}}/\|\mathbf{F}_{\text{R-ZFBF}}\|_{F}$, $\mathbf{W}_{\text{ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{ZFBF}}/\|\mathbf{F}_{\text{ZFBF}}\|_{F}$, and $\mathbf{W}_{\text{MBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{MBF}}/\|\mathbf{F}_{\text{MBF}}\|_{F}$, where the normalization factor has been included to ensure that the transmit power is the same as with digital beamforming.
In Fig. 5, we report the sum rate versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ achieved by digital and MiLAC-aided beamforming, where $N_{T}=N_{R}=4$ and the channels are independent and identically distributed (i.i.d.) Rayleigh distributed with unit path gain. We observe that R-ZFBF is the best precoder, which boils down to ZFBF at high SNR and to MBF at low SNR, for both digital and MiLAC-aided beamforming, as expected from MIMO theory [35, Chapter 12]. MiLAC-aided beamforming performs exactly as digital beamforming when the MiLAC is reconfigured based on the arbitrary beamforming matrix computed digitally, denoted as MiLAC (arbit.) in Fig. 5. Besides, MiLAC-aided beamforming suffers a slight performance degradation at high SNR when the LMMSE-inspired beamforming matrix is computed in the analog domain, referred to as MiLAC (LMMSE) in Fig. 5. This occurs because, in this case, MiLAC-aided beamforming cannot perform uniform power allocation, which is optimal at high SNR. At the cost of this slight performance degradation, computing the LMMSE-inspired beamforming matrices in the analog domain substantially reduces the required computational complexity, as it will be analyzed in the following.
Figure 7: Computational complexity per coherence block with length $\tau=100$ symbols of digital beamforming and MiLAC-aided beamforming
To compare MiLAC-aided beamforming and digital beamforming at the receiver, we consider a single-user MIMO system with an $N_{T}$-antenna transmitter and an $N_{R}$-antenna receiver, where the receiver combines the received signal through MMSE, ZF, or MF, as introduced in Section IV-C. In this case, the LMMSE-inspired combining matrices $\mathbf{G}_{\text{MMSE}}$, $\mathbf{G}_{\text{ZF}}$, and $\mathbf{G}_{\text{MF}}$ defined in Section IV-C for MiLAC-aided beamforming are also the optimal expressions of the MMSE, ZF, and MF receivers realized with digital beamforming.
In Fig. 6, we report the BER versus the SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$ obtained with digital and MiLAC-aided beamforming with $N_{T}=N_{R}=4$ and quadrature phase shift keying (QPSK) modulation. Since MiLAC-aided beamforming can implement the same combining matrices as digital beamforming, the two beamforming strategies achieve exactly the same BER.
(a) $\text{SNR}=0$ dB
(b) $\text{SNR}=10$ dB
(c) $\text{SNR}=20$ dB
Figure 8: Sum rate versus computational complexity of R-ZFBF performed in a $N_{R}\times N_{T}$ system with digital and MiLAC-aided beamforming, for three different levels of SNR.
VI-B Computational Complexity Evaluation
To quantify the benefits of MiLAC-aided beamforming in terms of computational complexity, we consider a MIMO system with $N_{T}$ transmitting antennas and $N_{R}$ receiving antennas, with $N_{T}=N_{R}$, where MiLAC-aided beamforming is used at the transmitter to perform R-ZFBF, ZFBF, and MBF, at the receiver for MMSE, ZF, and MF, and for DFT, as discussed in Section IV-B, IV-C, and IV-D. In the considered system, the beamforming matrix is reconfigured per coherence block, where each coherence block includes $\tau$ symbol transmissions. We define the computational complexity as the number of real operations per coherence block, as analyzed in the following.
  • Zero-forcing. For R-ZFBF and ZFBF at the transmitter, or MMSE and ZF at the receiver, MiLAC-aided beamforming requires $6N_{R}^{2}$ operations per coherence block, as specified in Remarks 2 and 3, and no operation is required on a per-symbol basis. Conversely, with digital beamforming, these strategies require $8(N_{R}^{3}+N_{R}^{3}/3)$ operations to design the beamforming matrix at each coherence block (see Remarks 2 and 3). In addition, a matrix-vector product ($\mathbf{W}\mathbf{s}$ or $\mathbf{G}\mathbf{y}$) is performed on a per-symbol basis, requiring $8N_{R}^{2}$ real operations per symbol time and making the complexity per coherence block of digital beamforming $8(N_{R}^{3}+N_{R}^{3}/3)+8N_{R}^{2}\tau$.
  • Matched filtering. For MBF at the transmitter, or MF at the receiver, MiLAC-aided beamforming requires $4N_{R}^{2}$ operations per coherence block, as given in Remarks 2 and 3, with no operation on a per-symbol basis. Besides, performing these strategies with digital beamforming requires $8N_{R}^{2}$ real operations per symbol time to perform the matrix-vector product $\mathbf{W}\mathbf{s}$ or $\mathbf{G}\mathbf{y}$, while no computation is needed to design the beamforming matrix, making the complexity per coherence block $8N_{R}^{2}\tau$.
  • DFT. Performing the DFT with MiLAC-aided beamforming requires no operation, since the MiLAC is not reconfigured at run-time (see Remark 4). Conversely, computing digitally the DFT requires $34/9N_{R}\log_{2}(N_{R})$ operations per symbol time [37], giving $34/9N_{R}\log_{2}(N_{R})\tau$ operations per coherence block.
In Fig. 7, we show the computational complexity of digital and MiLAC-aided beamforming, as a function of the number of antennas $N_{T}=N_{R}$, by fixing $\tau=100$ symbols per coherence block. For the cases zero-forcing and matched filtering, MiLAC-aided beamforming demonstrates up to $1.5\times 10^{4}$ and $2.0\times 10^{2}$ times lower computational complexity compared to digital beamforming, respectively, when the number of antennas is $N_{R}=8192$. This significant reduction comes because of two reasons. First, MiLAC-aided beamforming does not require the computation of the product $\mathbf{W}\mathbf{s}$, or $\mathbf{G}\mathbf{y}$, on a per-symbol basis, which has significant complexity given the high value of $\tau$ [3]. Second, for zero-forcing, the complexity of MiLAC-aided beamforming scales with $\mathcal{O}(N_{R}^{2})$ instead of $\mathcal{O}(N_{R}^{3})$ since the expensive computation of the beamforming matrix is not required. Besides, in the case of DFT, MiLAC-aided beamforming does not require any computation, saving up to $4.0\times 10^{7}$ operations per coherence block when $N_{R}=8192$. Given this significant reduction in computational complexity, we expect a proportional reduction in central processing unit (CPU) energy consumption and processing time.
VI-C Relationship Between Performance and Computational Complexity
We have shown that MiLAC-aided beamforming can achieve the same performance as digital beamforming in Figs. 5 and 6, and we have quantified its gain in terms of computational complexity in Fig. 7. We now investigate how the achieved performance and the required computational complexity are related in digital and MiLAC-aided beamforming. To this end, we consider multi-user MISO system with an $N_{T}$-antenna transmitter serving $N_{R}$ single-antenna receivers through R-ZFBF. For digital beamforming, we consider the R-ZFBF beamforming matrix to have uniform power allocation, as in Section VI-A. For MiLAC-aided beamforming, we consider the LMMSE-inspired design discussed in Section IV-B, where the computation of the R-ZFBF beamforming matrix is done in the analog domain.
In Fig. 8, we report the achieved sum rate and required computational complexity for MiLAC-aided and digital R-ZFBF beamforming in the case of different values of $N_{R}$ and $N_{T}$, where $N_{R},N_{T}\in\{256,1024,4096\}$ and $N_{R}\leq N_{T}$, and SNR, where $\text{SNR}\in\{0,10,20\}$ dB. We make the following three observations. First, the sum rate increases with the number of transmit antennas $N_{T}$, the number of receivers $N_{R}$, and the SNR, for both MiLAC-aided and digital beamforming. Besides, the computational complexity increases with the number of transmit antennas $N_{T}$ and receivers $N_{R}$, for both MiLAC-aided and digital beamforming, while is independent of the SNR. Second, with a high number of transmit antennas, i.e., $N_{T}\geq 256$, the performance achieved by a MiLAC computing the R-ZFBF matrix in the analog domain as per Section IV-B is the same as with digital beamforming with uniform power allocation. Thus, we can fully exploit the computational benefits of MiLAC-aided beamforming without sacrificing performance. Third, MiLAC-aided beamforming offers the same performance as digital beamforming for all the considered configurations of $N_{R}$, $N_{T}$, and SNR, with significant gains in computational complexity. Under a different perspective, MiLAC-aided beamforming can significantly improve the sum rate over digital beamforming given the same computational complexity. For example, MiLAC-aided beamforming can operate R-ZFBF in a system with $N_{R}=N_{T}=4096$ requiring approximately the same complexity as digital R-ZFBF in a system with $N_{R}=N_{T}=256$.
VII Conclusion
In Part II of this paper, we demonstrate the application of MiLAC to wireless communications, highlighting its potential to support communications with thousands of antennas by enabling analog-domain computations and beamforming. We first derive five special cases of the LMMSE estimator that have a practical interest in wireless communications, such as the LS and MF, and illustrate how they can be computed by a MiLAC in the analog domain. Following the capability of a MiLAC to compute such special cases, we introduce MiLAC-aided beamforming as a novel beamforming strategy enabled by a MiLAC. Compared to state-of-the-art beamforming strategies, MiLAC-aided beamforming offers the following five unique benefits in terms of performance, hardware complexity, and computational complexity. First, it has the same flexibility as digital beamforming, allowing us to achieve maximum performance. Second, it requires the minimum number of costly RF chains, i.e., equal to the number of transmitted symbols. Third, it can be realized with low-resolution ADCs/DACs since the symbols are precoded fully in the analog domain. Fourth, it does not perform any operation on a per-symbol basis, since it does not require precoding in the digital domain. Fifth, it can offer a significant computational complexity reduction in the case of ZF operations at the transmitter or at the receiver, which can be performed with complexity growing with the square of the antenna number, rather than the cube. Numerical results support the theoretical insights, showing that MiLAC-aided beamforming can offer a remarkable reduction in computational complexity. For example, ZF can be achieved with a complexity reduction of $1.5\times 10^{4}$ times.